OpenStudy (anonymous):

Fill in the missing value. 2/15 / [ ] = 2/3 Does anyone know what fraction is needed to be put in there? If so, please tell me how you do it.

OpenStudy (anonymous):

2/15 / x = 2/3, so multiplying both sides of the eqn by x gives 2/15 = x * 2/3. From here, dividing the right side of the eqn by 2/3 sets x by itself, giving x = (2/15)/(2/3) = 1/5

OpenStudy (anonymous):

Ohhh thanks :D